K Company estimates that overhead costs for the next year will be $3,500,000 for indirect labor and $940,000 for factory utilities. The company uses direct labor hours as its overhead allocation base. If 100,000 direct labor hours are planned for this next year, what is the company's plantwide overhead rate?

Multiple Choice

  • $0.02 per direct labor hour.

  • $44.40 per direct labor hour.

  • $35.00 per direct labor hour.

  • $9.40 per direct labor hour.

  • $0.11 per direct labor hour.
Answer #1

Answer : $ 44.40 Per Direct Labour Hours

Total estimated overheads = 3,500,000+940,000 = 4,440,000

Total Allocation base = 100,000 Direct labor hours

Plant overhead rate = Estiamted overhead / Total allocation BAse

Plant overhead Rate = 4,440,000 /100,000 = $44.40 Per DLH